OpenStudy (anonymous):

Evaluate the following integrals by converting to polar coordinates. 1 sqrt(1-x^2) ∫ ∫ e^(x^2+y^2)dydx 0 0 Please explain step by step

OpenStudy (anonymous):

a) Conversion to polar coordinates: Let I=∫11/2√∫x1−x2√1x2+y2−−−−−−√dydx. If you draw a diagram of the situation you will see that I=∫π/40∫r=secθr=11rrdrdθ=∫π/40secθ−1 dθ =[12log∣∣∣1+sinθ1−sinθ∣∣∣−θ]π/40=log(1+2√)−π4. (b) Standard hyperbolic substitution: To evaluate the inner integral set y=xsinhu and noting that sinh−1u=log(u+1+u2−−−−−√) we have ∫x1−x2√1x2+y2−−−−−−√dy=⎡⎣log⎛⎝yx+1+y2x2−−−−−−√⎞⎠⎤⎦x1−x2√ =log(1+2√)+logx−log(1+1−x2−−−−−√).(1) Both the logs can be integrated by parts. The first is standard, ∫logx dx=xlogx−x+C and the second ∫log(1+1−x2−−−−−√) dx=xlog(1+1−x2−−−−−√)−x+sin−1x+C. You will need (do the integration) to note that 11−x2−−−−−√−1=x2(1−x2)+1−x2−−−−−√. And so upon integrating (1) between 1/2√ and 1 you obtain I=(1−12√)log(1+2√)+[xlogx−xlog(1+1−x2−−−−−√)−sin−1x]11/2√ =log(1+2√)−π4.

OpenStudy (anonymous):

its a hint

OpenStudy (turingtest):

converting to polar coordinates is done with the formulas\[x=r\cos\theta\]\[y=r\sin\theta\]and in polar coordinates \[dA=rdrd\theta\]

OpenStudy (turingtest):

the area suggested by the first bound is the regionwhich is a circle of radius \(r=1\) therefor \(0\le r\le1\) and \(0\le\theta\le2\pi\) the exponent changes thanks to our conversion as well:\[x^2+y^2=r^2\cos^2\theta+r^2\sin^2\theta=r^2(\sin^2\theta+\cos^2\theta)=r^2\]putting this all together we get our new integral for the area\[\large A=\int\int dA=\int_{0}^{2\pi}\int_{0}^{1}re^{r^2}drd\theta\]

OpenStudy (turingtest):

actually I guess the last integral should be\[\large \int\int f(x)dA=\int_{0}^{2\pi}\int_{0}^{1}re^{r^2}drd\theta\]since you are integrating a function over the area

OpenStudy (turingtest):

@Janex does this make sense? can you do this integral?

OpenStudy (anonymous):

it made sense but when i was integrating i get \[r (e-1) 2\pi\] not sure what i did wrong

OpenStudy (turingtest):

\[\large\int_{0}^{2\pi}\int_{0}^{1}r e^{r^2}drd\theta=\int_{0}^{2\pi}\frac12e^{r^2}|_{0}^{1}d\theta=\frac12\int_{0}^{2\pi}e-1d\theta=\pi(e-1)\]do you see your mistake now?